# Unfitted hybrid high-order methods stabilized by polynomial extension for elliptic interface problems

### v1

@@ Introduction
Let $u$ be the weak solution to (2). Let $\hat{v}_{h}\in\widehat{\mathcal{U}}_{h0}$ be arbitrary and set
$\boldsymbol{\delta}_{T,i}$ $:=\boldsymbol{G}_{T^{i}}^{k}(\hat{v}_{T}^{\mathcal{N}})+\delta_{i1}\boldsymbol{L}_{T^{1}}^{k}(\llbracket u\rrbracket_{\Gamma})-\nabla u_{i}|_{T^{i}},$ $\forall(T,i)\in\mathcal{P}_{h},$ (36a)
$\boldsymbol{\delta}_{T,S,i}$ :=$:=\big{\{}\boldsymbol{G}_{T^{i}}^{k}(\hat{v}_{T}^{\mathcal{N}})+\delta_{i1}\boldsymbol{L}_{T^{1}}^{k}(\llbracket u\rrbracket_{\Gamma})\big{\}}^{+}|_{S^{i}}-\nabla u_{i}|_{S^{i}},$ $\forall(T,i)\in\mathcal{P}_{h}^{\textup{OK}},\,\forall S\in\mathcal{N}_{i}^{-1}(T),$ (36b)
where, consistently with the notation introduced above, the superscript + applied to a polynomial in $\mathbb{P}^{k}(T^{i};\mathbb{R}^{d})$ indicates its extension to $T^{i}\cup\bigcup_{S\in\mathcal{N}_{i}^{-1}(T)}S^{i}$. Then, the following holds for all $\hat{w}_{h}\in\widehat{\mathcal{U}}_{h0}$,
@@
$+s_{h}^{\circ}(\hat{v}_{h},\hat{w}_{h})+s_{h}^{\mathcal{N}}(\hat{v}_{h},\hat{w}_{h})+\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\kappa_{1}h_{T}^{-1}(\llbracket v_{T}-u\rrbracket_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}.$ (37)
@@
$L_{1}:={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}-(\nabla{\cdot}(\kappa_{i}\nabla u_{i}),w_{T^{i}})_{T^{i}}+\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}((\kappa_{1}\nabla u_{1}-\kappa_{2}\nabla u_{2}){\cdot}\boldsymbol{n}_{\Gamma},w_{T^{2}})_{T^{\Gamma}},$
$L_{2}:={}$ $\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\kappa_{1}h_{T}^{-1}(\llbracket u\rrbracket_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}},\qquad L_{3}:=-\sum_{(T,1)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{1}(\boldsymbol{L}^{k}_{T^{1}}(\llbracket u\rrbracket_{\Gamma}),\boldsymbol{G}_{T^{1}}^{k}(\hat{w}_{T}^{\mathcal{N}}))_{T^{1}}.$
@@
$\sum_{(T,i)\in\mathcal{P}_{h}}-(\nabla{\cdot}(\kappa_{i}\nabla u_{i}),w_{T^{i}})_{T^{i}}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\big{\{}(\nabla u_{i},\nabla w_{T^{i}})_{T^{i}}-(\nabla u_{i}{\cdot}\boldsymbol{n}_{T},w_{T^{i}})_{(\partial T)^{i}}\big{\}}$
$-\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\big{\{}(\kappa_{1}\nabla u_{1}{\cdot}\boldsymbol{n}_{\Gamma},w_{T^{1}})_{T^{\Gamma}}-(\kappa_{2}\nabla u_{2}{\cdot}\boldsymbol{n}_{\Gamma},w_{T^{2}})_{T^{\Gamma}}\big{\}}.$
Since $\nabla u_{i}$ is single-valued on $(\partial T)^{i}\cap\Omega_{i}$ and $w_{(\partial T)^{i}}$ vanishes on the mesh boundary faces, we obtain
$\sum_{(T,i)\in\mathcal{P}_{h}}-(\nabla{\cdot}(\kappa_{i}\nabla u_{i}),w_{T^{i}})_{T^{i}}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\big{\{}(\nabla u_{i},\nabla w_{T^{i}})_{T^{i}}+(\nabla u_{i}{\cdot}\boldsymbol{n}_{T},w_{(\partial T)^{i}}-w_{T^{i}})_{(\partial T)^{i}}\big{\}}$
$-\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\big{\{}(\kappa_{1}\nabla u_{1}{\cdot}\boldsymbol{n}_{\Gamma},w_{T^{1}})_{T^{\Gamma}}-(\kappa_{2}\nabla u_{2}{\cdot}\boldsymbol{n}_{\Gamma},w_{T^{2}})_{T^{\Gamma}}\big{\}}.$
@@
$L_{1}=\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\Big{\{}(\nabla u_{i},\nabla w_{T^{i}})_{T^{i}}+(\nabla u_{i}{\cdot}\boldsymbol{n}_{T},w_{(\partial T)^{i}}-w_{T^{i}})_{(\partial T)^{i}}\Big{\}}-\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\kappa_{1}(\nabla u_{1}{\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}.$
@@
$L_{1}=\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\Big{\{}(\nabla u_{i},\nabla w_{T^{i}})_{T^{i}}+(\nabla u_{i}{\cdot}\boldsymbol{n}_{T},w_{(\partial T)^{i}}-w_{T^{i}})_{(\partial T)^{i}}-\delta_{i1}(\nabla u_{1}{\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}\Big{\}}.$
@@
$L_{1}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}(\nabla u_{i},\nabla w_{T^{i}})_{T^{i}}$
$+\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}\bigg{\{}(\nabla u_{i}{\cdot}\boldsymbol{n}_{T},w_{(\partial T)^{i}}-w_{T^{i}})_{(\partial T)^{i}}-\delta_{i1}(\nabla u_{1}{\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}$
$+\sum_{S\in\mathcal{N}_{i}^{-1}(T)}\big{\{}(\nabla u_{i}{\cdot}\boldsymbol{n}_{S},w_{(\partial S)^{i}}-w_{S^{i}})_{(\partial S)^{i}}-\delta_{i1}(\nabla u_{1}{\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{S}\rrbracket_{\Gamma})_{S^{\Gamma}}\big{\}}\bigg{\}}.$
(2) For all $\hat{v}_{h}\in\widehat{\mathcal{U}}_{h0}$, setting $\boldsymbol{\gamma}_{T,i}:=\boldsymbol{G}_{T^{i}}^{k}(\hat{v}_{T}^{\mathcal{N}})+\delta_{i1}\boldsymbol{L}_{T^{1}}^{k}(\llbracket u\rrbracket_{\Gamma})$ for all $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$ and $\boldsymbol{\gamma}_{T,i}:=\boldsymbol{G}_{T^{i}}^{k}(\hat{v}_{T}^{\mathcal{N}})$ for all $(T,i)\in\mathcal{P}_{h}^{\textup{KO}}$, and using the definition (34) of the bilinear form $b_{h}$, we infer that
$b_{h}(\hat{v}_{h},\hat{w}_{h})-L_{3}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}(\boldsymbol{G}_{T^{i}}^{k}(\hat{v}_{T}^{\mathcal{N}}),\boldsymbol{G}_{T^{i}}^{k}(\hat{w}_{T}^{\mathcal{N}}))_{T^{i}}+\sum_{(T,1)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{1}(\boldsymbol{L}^{k}_{T^{1}}(\llbracket u\rrbracket_{\Gamma}),\boldsymbol{G}_{T^{1}}^{k}(\hat{w}_{T}^{\mathcal{N}}))_{T^{1}}$
[…]
@@
In this section, we present the main tools to perform the error analysis. The proofs are postponed to Section 6. For all $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$, the superscript + is used here to indicate the extension of a polynomial originally defined on $T^{i}$ to $\Delta(T)$ (recall that $T\cup\bigcup_{S\in\mathcal{N}_{i}^{-1}(T)}S\subset\Delta(T)$).
@@
where $\Pi^{k+1}_{T}$ denotes the $L^{2}$-orthogonal projection onto $\mathbb{P}^{k+1}(T;\mathbb{R})$. (Notice that the operator $I^{k+1}_{T^{i}}$ depends on the Sobolev index $s$, but this dependency is not tracked to simplify the notation; notice also that the extension operator is not needed if $T$ is an uncut cell.) Let us set
@@
$+h_{S}\|\nabla(v_{i}-I_{T^{i}}^{k+1}(v_{i})^{+})\|_{S^{i}}+h_{S}^{\frac{3}{2}}\|\nabla(v_{i}-I_{T^{i}}^{k+1}(v_{i})^{+})\|_{(\partial S)^{i}}\Big{\}}.$ (40)
@@
$\sum_{S\in\{T\}\cup\mathcal{N}_{i}^{-1}(T)}h_{S}^{\frac{1}{2}}\|\llbracket v-I_{T}^{k+1}(v)^{+}\rrbracket_{\Gamma}\|_{S^{\Gamma}}$ $\lesssim h_{T}^{s}\sum_{i\in\{1,2\}}|E_{i}^{s}(v_{i})|_{H^{s}(\Delta_{2}(T))},$ (41b)
@@
$+\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\sum_{S\in\mathcal{N}_{i}^{-1}(T)}\kappa_{i}h_{S}^{-1}\|v_{S^{i}}-v_{T^{i}}^{+}\|_{(\partial S)^{i}}^{2},$ (44)
where (we recall that) $v_{T^{i}}^{+}$ denotes the extension of the polynomial $v_{T^{i}}\in\mathbb{P}^{k+1}(T^{i};\mathbb{R})$ to $T^{i}\cup\bigcup_{S\in\mathcal{N}_{i}^{-1}(T)}S^{i}$. It is straightforward to verify that (42) defines a norm on $\widehat{\mathcal{U}}_{h0}$. Indeed, if $\hat{v}_{h}\in\widehat{\mathcal{U}}_{h0}$ satisfies $\|\hat{v}_{h}\|_{h0}=0$, then all the cell unknowns and all the face unknowns are constant and take the same value inside each sub-domain $\Omega_{i}$ and globally in the domain $\Omega$. Since the face unknowns on the boundary $\partial\Omega$ vanish, we conclude that all the components of $\hat{v}_{h}$ are zero.
@@
Squaring this inequality, multiplying by $\kappa_{i}h_{S}^{-1}$, summing over $S\in\mathcal{N}_{i}^{-1}(T)$, combining with (47), and finally summing over $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$, we infer that

### v2

@@ Introduction
Let $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}$} be the weak solution to (2). Let $\hat{v}_{h}\in\widehat{\mathcal{U}}_{h0}$ be arbitrary and set
$\boldsymbol{\delta}_{T,i}$ $:=\boldsymbol{G}_{T^{i}}^{k}(\hat{v}_{T}^{\mathcal{N}})+\delta_{i1}\boldsymbol{L}_{T^{1}}^{k}(\llbracket{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\rrbracket_{\Gamma})-\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}|_{T^{i}},$ $\forall(T,i)\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}_{h}^{\textup{OK}}},$ (36a)
$\boldsymbol{\delta}_{T,S,i}$ $:=\big\{\boldsymbol{G}_{T^{i}}^{k}(\hat{v}_{T}^{\mathcal{N}})+\delta_{i1}\boldsymbol{L}_{T^{1}}^{k}(\llbracket{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\rrbracket_{\Gamma})\big\}^{+}|_{S^{i}}-\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}|_{S^{i}},$ $\forall(T,i)\in\mathcal{P}_{h}^{\textup{OK}},\,\forall S\in\mathcal{N}_{i}^{-1}(T),$ (36b)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\delta}_{T,i}}$ $:={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{G}_{T^{i}}^{k}(\hat{v}_{T}^{\mathcal{N}})-\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}|_{T^{i}}=\nabla v_{T^{i}}-\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}|_{T^{i}},}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\forall(T,i)\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}_{h}^{\textup{KO}}}}.$ (36c)
Then, the following holds for all $\hat{w}_{h}\in\widehat{\mathcal{U}}_{h0}$,
@@
$+s_{h}^{\circ}(\hat{v}_{h},\hat{w}_{h})+s_{h}^{\mathcal{N}}(\hat{v}_{h},\hat{w}_{h})+\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\kappa_{1}h_{T}^{-1}(\llbracket v_{T}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\rrbracket_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}.$ (37)
@@
$L_{1}:={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}-(\nabla{\cdot}(\kappa_{i}\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}),w_{T^{i}})_{T^{i}}+\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}((\kappa_{1}\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{1}}-\kappa_{2}\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{2}}){\cdot}\boldsymbol{n}_{\Gamma},w_{T^{2}})_{T^{\Gamma}},$
$L_{2}:={}$ $\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\kappa_{1}h_{T}^{-1}(\llbracket{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\rrbracket_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}},\qquad L_{3}:=-\sum_{(T,1)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{1}(\boldsymbol{L}^{k}_{T^{1}}(\llbracket{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\rrbracket_{\Gamma}),\boldsymbol{G}_{T^{1}}^{k}(\hat{w}_{T}^{\mathcal{N}}))_{T^{1}}.$
@@
$\sum_{(T,i)\in\mathcal{P}_{h}}-(\nabla{\cdot}(\kappa_{i}\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}),w_{T^{i}})_{T^{i}}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\big\{(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}},\nabla w_{T^{i}})_{T^{i}}-(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}{\cdot}\boldsymbol{n}_{T},w_{T^{i}})_{(\partial T)^{i}}\big\}$
$-\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\big\{(\kappa_{1}\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{1}}{\cdot}\boldsymbol{n}_{\Gamma},w_{T^{1}})_{T^{\Gamma}}-(\kappa_{2}\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{2}}{\cdot}\boldsymbol{n}_{\Gamma},w_{T^{2}})_{T^{\Gamma}}\big\}.$
Since $\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}$ is single-valued on $(\partial T)^{i}\cap\Omega_{i}$ and $w_{(\partial T)^{i}}$ vanishes on the mesh boundary faces, we obtain
$\sum_{(T,i)\in\mathcal{P}_{h}}-(\nabla{\cdot}(\kappa_{i}\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}),w_{T^{i}})_{T^{i}}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\big\{(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}},\nabla w_{T^{i}})_{T^{i}}+(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}{\cdot}\boldsymbol{n}_{T},w_{(\partial T)^{i}}-w_{T^{i}})_{(\partial T)^{i}}\big\}$
$-\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\big\{(\kappa_{1}\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{1}}{\cdot}\boldsymbol{n}_{\Gamma},w_{T^{1}})_{T^{\Gamma}}-(\kappa_{2}\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{2}}{\cdot}\boldsymbol{n}_{\Gamma},w_{T^{2}})_{T^{\Gamma}}\big\}.$
@@
$L_{1}=\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\Big\{(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}},\nabla w_{T^{i}})_{T^{i}}+(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}{\cdot}\boldsymbol{n}_{T},w_{(\partial T)^{i}}-w_{T^{i}})_{(\partial T)^{i}}\Big\}-\sum_{T\in\mathcal{T}_{h}^{\textup{cut}}}\kappa_{1}(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{1}}{\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}.$
@@
$L_{1}=\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}\Big\{(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}},\nabla w_{T^{i}})_{T^{i}}+(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}{\cdot}\boldsymbol{n}_{T},w_{(\partial T)^{i}}-w_{T^{i}})_{(\partial T)^{i}}-\delta_{i1}(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{1}}{\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}\Big\}.$
@@
$L_{1}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}},\nabla w_{T^{i}})_{T^{i}}$
$+\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}\bigg\{(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}{\cdot}\boldsymbol{n}_{T},w_{(\partial T)^{i}}-w_{T^{i}})_{(\partial T)^{i}}-\delta_{i1}(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{1}}{\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{T}\rrbracket_{\Gamma})_{T^{\Gamma}}$
$+\sum_{S\in\mathcal{N}_{i}^{-1}(T)}\big\{(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{i}}{\cdot}\boldsymbol{n}_{S},w_{(\partial S)^{i}}-w_{S^{i}})_{(\partial S)^{i}}-\delta_{i1}(\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}_{1}}{\cdot}\boldsymbol{n}_{\Gamma},\llbracket w_{S}\rrbracket_{\Gamma})_{S^{\Gamma}}\big\}\bigg\}.$
(2) For all $\hat{v}_{h}\in\widehat{\mathcal{U}}_{h0}$, setting $\boldsymbol{\gamma}_{T,i}:=\boldsymbol{G}_{T^{i}}^{k}(\hat{v}_{T}^{\mathcal{N}})+\delta_{i1}\boldsymbol{L}_{T^{1}}^{k}(\llbracket{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\rrbracket_{\Gamma})$ for all $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$ and $\boldsymbol{\gamma}_{T,i}:=\boldsymbol{G}_{T^{i}}^{k}(\hat{v}_{T}^{\mathcal{N}}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\nabla v_{T^{i}}}$ for all $(T,i)\in\mathcal{P}_{h}^{\textup{KO}}$, and using the definition (34) of the bilinear form $b_{h}$ together with the definition of $L_{3}$ given above, we infer that
$b_{h}(\hat{v}_{h},\hat{w}_{h})-L_{3}={}$ $\sum_{(T,i)\in\mathcal{P}_{h}}\kappa_{i}(\boldsymbol{G}_{T^{i}}^{k}(\hat{v}_{T}^{\mathcal{N}}),\boldsymbol{G}_{T^{i}}^{k}(\hat{w}_{T}^{\mathcal{N}}))_{T^{i}}+\sum_{(T,1)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{1}(\boldsymbol{L}^{k}_{T^{1}}(\llbracket{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\rrbracket_{\Gamma}),\boldsymbol{G}_{T^{1}}^{k}(\hat{w}_{T}^{\mathcal{N}}))_{T^{1}}$
$={}$ $\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\kappa_{i}(\boldsymbol{\gamma}_{T,i},\boldsymbol{G}_{T^{i}}^{k}(\hat{w}_{T}^{\mathcal{N}}))_{T^{i}}+\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{KO}}}\kappa_{i}(\boldsymbol{\gamma}_{T,i},\boldsymbol{G}_{T^{i}}^{k}(\hat{w}_{T}^{\mathcal{N}}))_{T^{i}}.$
@@
In this section, we present the main tools to perform the error analysis. The proofs are postponed to Section 6. For all $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$, recall that the superscript + is used here to indicate the extension of a polynomial originally defined on $T^{i}$ to $\Delta(T)$ (observe that $T\cup\bigcup_{S\in\mathcal{N}_{i}^{-1}(T)}S\subset\Delta(T)$).
@@
where $\Pi^{k+1}_{T}$ denotes the $L^{2}$-orthogonal projection onto $\mathbb{P}^{k+1}(T;\mathbb{R})$. (Notice that the operator $I^{k+1}_{T^{i}}$ depends on the Sobolev index $s$, but this dependency is not tracked to simplify the notation; notice also that the extension operator is not needed if $T$ is an uncut cell.) Let us set, for all $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$,
@@
$\hskip 49.79231pt+h_{S}\|\nabla(v_{i}-I_{T^{i}}^{k+1}(v_{i})^{+})\|_{S^{i}}+h_{S}^{\frac{3}{2}}\|\nabla(v_{i}-I_{T^{i}}^{k+1}(v_{i})^{+})\|_{(\partial S)^{i}}\Big\},$ (40a)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(v)}:={}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{S\in\{T\}\cup\mathcal{N}_{i}^{-1}(T)}h_{S}^{\frac{1}{2}}\|\llbracket v-I_{T}^{k+1}(v)^{+}\rrbracket_{\Gamma}\|_{S^{\Gamma}}},$ (40b)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon^{\Gamma}_{T,i}(v)}$ $\lesssim h_{T}^{s}\sum_{i\in\{1,2\}}|E_{i}^{s}(v_{i})|_{H^{s}(\Delta_{2}(T))}.$ (41b)
@@
$+\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\sum_{S\in\mathcal{N}_{i}^{-1}(T)}\eta_{\mathcal{N}}\kappa_{i}h_{T}^{-2}\|v_{S^{i}}-v_{T^{i}}^{+}\|_{T^{i}}^{2}.$ (44)
It is straightforward to verify that (42) defines a norm on $\widehat{\mathcal{U}}_{h0}$. Indeed, if $\hat{v}_{h}\in\widehat{\mathcal{U}}_{h0}$ satisfies $\|\hat{v}_{h}\|_{h0}=0$, then all the cell unknowns and all the face unknowns are constant and take the same value inside each sub-domain $\Omega_{i}$ and globally in the domain $\Omega$. Since the face unknowns on the boundary $\partial\Omega$ vanish, we conclude that all the components of $\hat{v}_{h}$ are zero.
@@
Invoking the Cauchy–Schwarz inequality and the discrete trace inequality (38a), we obtain
@@
The first term on the right-hand side is bounded using the $L^{2}$-stability of $\Pi_{(\partial S)^{i}}^{k}$ and (47). The second term is controlled by means of the discrete trace inequality (38a) and the definition (30) of the stabilization bilinear from $s_{h}{{}^{\mathcal{N}}}$. Squaring the resulting inequality, multiplying by $\kappa_{i}h_{S}^{-1}$, summing over $S\in\mathcal{N}_{i}^{-1}(T)$, and finally summing over $(T,i)\in\mathcal{P}_{h}^{\textup{OK}}$, we obtain